. Howton & Howton Worldwide (HHW) is planning its operations for the coming year, and the CEO wants you to forecast the firm's additional funds needed (AFN). The firm is operating at full capacity. Data for use in the forecast are shown below. However, the CEO is concerned about the impact of a change in the payout ratio from the 10% that was used in the past to 50%, which the firm's investment bankers have recommended. Based on the AFN equation, by how much would the AFN for the coming year change if HHW increased the payout from 10% to the new and higher level? All dollars are in millions.
Last year’s sales = S0 $300.0 Last year’s accounts payable $50.0
Sales growth rate = g 40% Last year’s notes payable $15.0
Last year’s total assets = A0* $500.0 Last year’s accruals $20.0
Last year’s profit margin = PM 20.0% Initial payout ratio 10.0%
1
Expert's answer
2012-02-23T09:38:13-0500
Answer:


the increase of AFN = $33.6 mln


Explanation:

AFN = projected increase in assets – spontaneous increase in liabilities – increase in retained earnings

The company is at full capacity, so assets must grow at the same rate as projected sales: $500*1.4=$700, projected increase in assets = $700 - $500 =
$200

Total sales = $300 *1.4 = $420

spontaneous increase in liabilities = X, 20/500 = X/700 => X = 28

For payout ratio = 10%:
Increase in Retained earnings = Net Income = 420 X 20% = 84, Dividend = 10% =
84X10%=8.4.
Increase in retained earnings = 84-8.4 = 75.6
AFN = $200 – $28 – $75.6 = $96.4 million

For payout ratio = 50%:
Increase in Retained earnings = Net Income = $420 * 20% = $84, Dividend = 50% = $84 * 50%=$42.
Increase in retained earnings = $84 - $42 = $42
AFN = $200 – $28 – $42 = $130 million

AFN change = $130 - $96.4 = $33.6